OpenStudy (anonymous):

open study evaluate the expression (4x^3)^2 for x = 2

OpenStudy (anonymous):

262144

OpenStudy (anonymous):

Thats wrong

OpenStudy (rushwr):

\[(4x ^{3})^{2}\] This is it right?

OpenStudy (javalos01):

yea it is

OpenStudy (javalos01):

i agree

OpenStudy (rushwr):

\[(4x ^{3})^{2} = 4x ^{6} = 4 * 2^{6} = 4* 64 = \]

OpenStudy (rushwr):

Hope u got it @Jesus6757

OpenStudy (rushwr):

The answer would be 256 ! :)

OpenStudy (javalos01):

good job @Rushwr